# Rethinking Open-Vocabulary Segmentation of Radiance Fields in 3D Space

### v1

@@ 4.2 Computational Time
Table 4 shows the computational times of the LERF waldo_kitchen scene with a single RTX A5000. We report the training time and rendering time for both ours and the competitors. LangSplat and LEGaussians train an autoencoder for each scene and require decoding during rendering. Note that LangSplat trains 3DGS from scratch, optimizes the language-embedded 3DGS per scale, and requires additional post-processing. Ours-3DGS takes 40 minutes to train Ours-NeRF, 5 minutes to optimize 3DGS, and 22 milliseconds to query and transfer language field into 3DGS.
Ours-3DGS achieves real-time rendering of language embeddings for the first time, 27 faster than LEGaussians. Note that LangSplat heavily depends on post-processing as shown in Appendix B.2.
@@ 4.3 Ablation Study
We demonstrate the necessity of our point-wise semantic loss. Figure 9 qualitatively shows improvements due to point-wise semantic losscompared to 2D supervision with Eq. (1). point-wise semantic lossremoves meaningless floaters related to depth ambiguity. Table 6 shows that leveraging point-wise semantic loss instead of Eq. (1) makes better understanding in 3D space, improving the F1-score (+0.0995). Furthermore, point-wise semantic loss also indirectly improves 2D segmentation.
@@
Our scale-free language embedding field from SAM-segmented objects produce greatly improved accuracy with clear object boundaries compared to multi-scale embedding field. Figure 9 shows that scale-free embedding helps covering the head of the toy elephant which was lost with multi-scale embedding. In Table 6, the quantitative performance of the 3D and 2D segmentation greatly improves using scale-free embedding.
| Method | 3D Segmentation | 2D Segmentation | |
| F1-score | mIoU | mAP | |
| Ours full | 0.1532 | 46.37 | 45.86 |
| w/o point-wise semantic loss | 0.0537 | 44.50 | 44.05 |
| w/o scale-free embedding | 0.1114 | 30.03 | 28.85 |
@@
We demonstrate the necessity of initializing 3DGS with the extracted point clouds from our pre-trained NeRF. As shown in Table 5, training 3DGS with SfM (Schönberger and Frahm 2016) initialization with the densification and pruning leads to performance degradation in both 3D and 2D segmentation. This indicates that aligning geometry and semantics is essential for transferring the language field to 3DGS. Moreover, NeRF initialization produces fewer Gaussians (0.33) which lead to less training time (0.38) and faster rendering speed (2.7).
@@ 5 Conclusion
We revisit the current literature on 3D understanding of NeRF and 3DGS and revise the problem setting. We reformulate the task to produce 3D segmented volumes instead of 2D rendered masks and propose a 3D evaluation protocol. We achieve state-of-the-art segmentation accuracy in both 3D and 2D. Moreover, by transferring the language field to 3DGS, we enable the first real-time rendering speed. We hope this paper drives forward a better 3D understanding of radiance fields by reconsidering the problem set.
@@ A.2 Details of Datasets
We use three datasets for evaluation: Replica dataset for 3D segmentation, and LERF and 3D-OVS datasets for 2D segmentation. All of these datasets are widely used to evaluate the segmentation performance of radiance fields.
@@
3D-OVS dataset is a collection of real-world scenes, text queries, and their ground truth pixel-level labels. The dataset consists of 10 scenes, each with 20-30 viewpoints, including five evaluation viewpoints. Several scenes in 3D-OVS dataset show poor reconstruction quality from the evaluation view shown in Figure S1. Therefore, we report the performance only for the scenes with successful reconstruction: (bed, bench, lawn, and office_desk).
@@ Appendix B Details of Previous Language Embed 3DGS
B.1 Dimensions of Embeddings on 3DGS
Rendering 512-dimensional CLIP embeddings requires 522 KB of shared memory per streaming multiprocessor (SM). Specifically, 3DGS (Kerbl et al. 2023) requires 10 KB of shared memory and needs an additional 1 KB/SM for each increase in the dimension of the language feature. 3DGS allocates static shared memory, and CUDA limits this allocation to 48 KB for architectural compatibility. Even with a dynamic allocation that enables allocation above the 48 KB limit, 522 KB of shared memory exceeds the maximum shared memory of current top GPUs (e.g., the H100 has a maximum shared memory of 228 KB).
Consequently, existing methods either 1) compress high-dimensional features or 2) allocate variables for the backward pass to global memory instead of shared memory. Each approach has trade-offs: 1) compressing features reduces segmentation performance but accelerates training, while 2) allocating to global memory avoids performance degradation but slows down training. Compressing the features produces poor mIoU (-7.30) but speeds up the training time 47, as shown in Table S2. However, Ours-3DGS achieves fast training without feature compression.
| Method | LERF Dataset | ||
| mIoU | mAP | Training Time | |
@@ B.2 Post-process of LangSplat
We show the performance difference with and without post-processing in LangSplat222https://github.com/minghanqin/LangSplat, as shown in Figure S2. Without post-processing, it achieves a fast rendering speed of 55 fps (550x) but produces a poor mIoU (-5.5) as shown in Table S3. Note that Ours-3DGS achieve faster rendering speed even compared to LangSplat without post-processing, as shown in Table 4.
@@ Appendix C More Results
We report the quantitative results of each scene on Replica dataset, LERF dataset, and 3D-OVS dataset, in Table S4-Table S6.
| Method | office_0 | office_1 | office_2 | office_3 |
| F1-score | F1-score | F1-score | F1-score | |
| LERF | 0.0774 | 0.0191 | 0.0644 | 0.0621 |
| OpenNeRF | 0.0486 | 0.0169 | 0.0318 | 0.0356 |
| Ours-NeRF | 0.1434 | 0.0434 | 0.1493 | 0.1255 |
| LEGaussians | 0.0013 | 0.0043 | 0.0109 | 0.0031 |
| LangSplat | 0.0004 | 0.0123 | 0.0044 | 0.0038 |
| Ours-3DGS | 0.1097 | 0.0406 | 0.1395 | 0.1181 |
| Method | office_4 | room_0 | room_1 | room_2 |
| F1-score | F1-score | F1-score | F1-score | |
| LERF | 0.0645 | 0.1124 | 0.1607 | 0.1158 |
| OpenNeRF | 0.0521 | 0.0180 | 0.0432 | 0.0423 |
| Ours-NeRF | 0.1722 | 0.1654 | 0.1934 | 0.2232 |
| LEGaussians | 0.0111 | 0.0080 | 0.0066 | 0.0085 |
| LangSplat | 0.0107 | 0.0139 | 0.0122 | 0.0118 |

### v3

@@ 4.2 Computational Time
Table 3 shows the computational times of the LERF waldo_kitchen scene with a single RTX A5000. We report the training time and rendering time for both ours and the competitors. LangSplat and LEGaussians train an autoencoder for each scene and require decoding during rendering. Note that LangSplat trains 3DGS from scratch, optimizes the language-embedded 3DGS per scale, heavily depends on post-processing (see Appendix). Ours-3DGS takes 40 minutes to train Ours-NeRF, 5 minutes to optimize 3DGS, and 22 milliseconds to query and transfer language field into 3DGS. Ours-3DGS achieves real-time rendering of language embeddings for the first time, 28 faster than LEGaussians.
Method | Replica | LERF | |
F1-score | mIoU | mAP | |
Ours full | 0.1532 | 46.37 | 45.86 |
w/o | 0.0537 | 44.50 | 44.05 |
w/o | 0.1114 | 30.03 | 28.85 |
@@ 4.3 Ablation Study
We demonstrate the necessity of our point-wise semantic loss. Figure 9 qualitatively shows improvements due to point-wise semantic loss compared to 2D supervision with Eq. (1). Point-wise semantic loss removes meaningless floaters related to depth ambiguity. Table 5 shows that leveraging point-wise semantic loss instead of Eq. (1) makes better understanding in 3D space, improving the F1-score (+0.0995). Furthermore, point-wise semantic loss also indirectly improves rendered 2D segmentation accuracy.
@@
Our scale-free language embedding field from SAM-segmented objects produce greatly improved accuracy with clear object boundaries compared to multi-scale embedding field. Figure 9 shows that scale-free embedding helps covering the head of the toy elephant which was lost with multi-scale embedding. In Table 5, the quantitative performance of the 3D and rendered 2D segmentation greatly improves using scale-free embedding.
@@
We demonstrate the necessity of initializing 3DGS with the extracted point clouds from our learned NeRF. As shown in Table 4, training 3DGS with SfM (Schönberger and Frahm 2016) initialization with the densification and pruning leads to performance degradation in both 3D and rendered 2D segmentation. This indicates that aligning geometry and semantics is essential for transferring the language field to 3DGS. Moreover, NeRF initialization produces fewer Gaussians (0.50) which lead to less training time (0.38) and faster rendering speed (1.70).
@@ 5 Conclusion
We revisit the current literature on 3D understanding of NeRF and 3DGS and revise the problem setting. We reformulate the task to produce 3D segmented volumes instead of rendered 2D masks and propose a 3D evaluation protocol. We achieve state-of-the-art segmentation accuracy in both 3D and rendered 2D by computing loss directly on 3D points. Moreover, we enable the first real-time rendering speed among open-vocabulary methods by transferring the learned language field to 3DGS. We hope this paper drives forward a better 3D understanding of radiance fields by reconsidering the problem set.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No. RS-2024-00439762, Developing Techniques for Analyzing and Assessing Vulnerabilities, and Tools for Confidentiality Evaluation in Generative AI Models)
@@ A.2 Details of Datasets
We use three datasets for evaluation: Replica dataset for 3D segmentation, and LERF and 3D-OVS datasets for rendered 2D segmentation. All of these datasets are widely used to evaluate the segmentation performance of radiance fields. We fix the thresholds for each dataset as detailed in Table 9 to evaluate the performance.
@@
3D-OVS dataset is a collection of real-world scenes, text queries, and their ground truth pixel-level labels. The dataset consists of 10 scenes, each with 20-30 viewpoints, including five evaluation viewpoints. Several scenes in 3D-OVS dataset show poor reconstruction quality from the evaluation view shown in Figure 5. Therefore, we report the performance only for the scenes with successful reconstruction: (bed, bench, lawn, and office_desk).
ScanNet dataset.
ScanNet dataset is a real-world dataset that includes RGB-D scans, meshes, and instance-level semantic labels. It is widely used to evaluate performance in 3D scene understanding tasks. In our case, we select one scene (scene0003) and provide the 3D segmentation accuracy, in Appendix C.
@@ Appendix B Details of Previous Language Embed 3DGS
B.1 Splatting High Dimensions Embeddings on 3DGS
To jointly train 3DGS with language features, a straightforward approach is to embed the 512-dimensional language feature into each Gaussian. However, this approach results in out-of-memory(OOM) issues. Splatting 512-dimensional language embeddings requires 522 KB of shared memory per streaming multiprocessor (SM). Specifically, 3DGS (Kerbl et al. 2023) requires 10 KB of shared memory and needs an additional 1 KB/SM for each increase in the dimension of the language feature. 3DGS allocates static shared memory, and CUDA limits this allocation to 48 KB for architectural compatibility. Even with a dynamic allocation that enables allocation above the 48 KB limit, 522 KB of shared memory exceeds the maximum shared memory of current top GPUs (e.g., the H100 has a maximum shared memory of 228 KB).
Consequently, existing methods address these challenges using the following strategies, each with its trade-offs:
Allocate rasterizer variables for the backward pass to global memory instead of shared memory (Zhou et al. 2023). While this avoids performance degradation, it significantly slows down training.
Use dynamic shared memory allocation with reduced tile size in the rasterizer (Ye et al. 2024). This method is faster than global memory allocation while using more shared memory, but it still slows down training due to the smaller tile size.
Render in chunks (Bhalgat et al. 2024). Although this accelerates training and prevents performance degradation, it drops fps (10 fps reported in the original paper).
Compressing the features produces poor mIoU (-7.30) but speeds up the training time 47, as shown in Table 11. However, Ours-3DGS achieves fast training without feature compression.
Method | mIoU | mAP | Training Time |
@@ B.2 Post-process of LangSplat
We show the performance difference with and without post-processing in LangSplat, as shown in Figure 6. Without post-processing, it achieves a fast rendering speed of 55 fps (550) 444The post-processing consists of a mean convolution filter on the relevancy map implemented with OpenCV (61.98ms) and a mean convolution filter on the binary mask implemented with nested for-loops (3249 ms). This is done for three scales, 10473 ms in total. but produces a poor mIoU (-5.5) as shown in Table 12. Note that Ours-3DGS achieve faster rendering speed even compared to LangSplat without post-processing, as shown in Table 3.
@@ Appendix C More Results
We mainly evaluate 3D segmentation accuracy on Replica datasets. To expand this, we also evaluate and compare on scene_0003 of ScanNet (Dai et al. 2017) dataset, in Table 13.
Additionally, we report the quantitative results of each scene on Replica dataset, LERF dataset, and 3D-OVS dataset, in Table 14-16.
Backbone | Method | ScanNet |
F1-score | ||
NeRF | LERF | 0.0333 |
OpenNeRF | 0.0663 | |
Ours-NeRF | 0.0849 | |
3DGS | LEGaussians | 0.0064 |
LangSplat | 0.0340 | |
Ours-3DGS | 0.0769 |
Backbone | Method | office_0 | office_1 | office_2 | office_3 |
F1-score | F1-score | F1-score | F1-score | ||
NeRF | LERF | 0.0774 | 0.0191 | 0.0644 | 0.0621 |
OpenNeRF | 0.0486 | 0.0169 | 0.0318 | 0.0356 | |
Ours-NeRF | 0.1434 | 0.0434 | 0.1493 | 0.1255 | |
3DGS | LEGaussians | 0.0013 | 0.0043 | 0.0109 | 0.0031 |
LangSplat | 0.0004 | 0.0123 | 0.0044 | 0.0038 | |
Ours-3DGS | 0.1097 | 0.0406 | 0.1395 | 0.1181 | |
Backbone | Method | office_4 | room_0 | room_1 | room_2 |
F1-score | F1-score | F1-score | F1-score | ||
NeRF | LERF | 0.0645 | 0.1124 | 0.1607 | 0.1158 |
OpenNeRF | 0.0521 | 0.0180 | 0.0432 | 0.0423 | |
Ours-NeRF | 0.1722 | 0.1654 | 0.1934 | 0.2232 | |
3DGS | LEGaussians | 0.0111 | 0.0080 | 0.0066 | 0.0085 |
LangSplat | 0.0107 | 0.0139 | 0.0122 | 0.0118 | |